# The self-interaction effects on the Kerr black hole superradiance and their observational implications

### v1

@@ II.3 Self-interacting scalar bosons
Axions, or axion-like particles (ALPs), are promising candidates for dark matter. The searches for ultralight scalars like axions to date have not adequately accounted for self-interactions among particles in the superradiance process. Self-interactions, that could significantly alter superradiance dynamics, are generically present for many scalar models, including axion-like dark matter candidates and string axions. Besides the QCD axion arising as the result of the Peccei-Quinn symmetry broken at a high energy scale $f_{a}$, more general ALPs can be present in particular theory and are typically predicted to arise in string theory, in which the symmetry-breaking scale $f_{a}$ can fall in a wide mass range [9, 36]. From a particle physics perspective, the simplest interaction of a scalar field is the quartic self-interaction.
@@
$V_{\rm free}(\phi)=\frac{1}{2}{m_{a}^{2}}\phi^{2}\,\,.$ (19)
Therefore one can have the Klein-Gordon equation
$(\nabla^{2}-m_{a}^{2})\phi=0$ (20)
for the free field. The metric convention here is (,+,+,+)(-,+,+,+)( - , + , + , + ).
Self-interacting ultralight bosons generally arising through non-perturbative theory at some high energy scale give an effective potential. Any type of potential of the scalar field $\phi$ can be expanded to the lowest order with the mass term and the quartic self-interaction term. The strength of the mass term in the potential is proportional to the square of scalar mass, and the quartic self-interaction term proportional to the dimensionless self-interaction parameter $\lambda$. That is,
@@
$\lambda\approx 10^{-82}\left(\frac{m_{a}}{10^{-13}\leavevmode\nobreak\ \mathrm{eV}}\right)^{2}\left(\frac{10^{19}\leavevmode\nobreak\ \mathrm{GeV}}{f_{a}}\right)^{2}.$ (24)
@@
To the lowest order of $\phi$, heuristically Eq. (25) can be considered as the new Klein-Gordon equation in the Kerr metric and the deviation comes from the modification of the scalar mass. By considering the $\phi^{2}$ part as the background field, we can investigate this mass modification. The main contribution stems from the dominating energy level of bound state which grows at the maximum speed and occupies much more than others. The particle occupation number of the fastest growing level $N$ is proportional to the square of its field amplitude ($N\propto|\phi|^{2}$). We separate this part individually and define $\Pi$ to quantify the contribution of bound state wave function except $N$, i.e. $N\Pi={\phi^{2}}/\left.{f_{a}^{2}}\right.$. The quantity $\Pi$ is only determined by the distribution of the wave function and the scale of $f_{a}$. We are going to evaluate this modification in a more realistic situation for the superradiant growth. The realistic approach is based on the fact that for the gravitational atom’s parameter space of interest the growth timescale of the scalar cloud is much longer than the oscillation timescale of the particles, ${\omega}^{-1}$ ($\approx m_{a}^{-1}$). Also, the cloud growth timescale is much slower than the timescale of the particles’ orbital motion in the vicinity of the BH. For these reasons, we can perform the time- and spatial averaging when investigating $\Pi$ within one period and calculating the scalar mass deviation. Using the similar notation as in Eq. (7) and Eq. (8), we can then calculate $\Pi$ with
@@
where $\Gamma_{\rm SR}$ is the growth rate of bosons, $t$ is the time and the bars on the symbols denote the average over spatial variables. In the non-relativistic limit, we already know the analytical expressions of radial and angular functions. The radial dependence of function, $R$, becomes radial hydrogen wave function, while the angular dependence of function, $S$, becomes the spherical harmonics. Based on the more thorough calculation, although the analytical results provide a good approximation, we carry out the estimation in the full Kerr background. We can investigate the mean values of the functions around the cloud region. Note that the radial and angular functions both obey the normalization condition. As for the occupation number at a specific time, we can estimate $N$ through the spin difference of the BH due to the conservation of angular momentum. When only considering the fastest growing level of bound state, the scalar cloud is dominated by the particles of single state and the extracted angular momentum from Kerr BH is stored in the same state. The particle number of the fastest growing state is
@@
where $\varDelta\chi\lesssim\mathcal{O}(0.1)$. Owing to the hierarchy of growth rates for different superradiant state, the newly-formed BH is always dominated by one bound state of scalar boson, and higher states would take much longer time to grow to the saturation. Thus, we consider almost all bound-state bosons occupy the same energy level here.
@@ II.4 The spectra of the gravitational atom
We can apply the modification on the recurrence relation Eq. (14) and obtain the new continued fraction equation for self-interacting bound state boson. While the deviation of boson mass induced by the self-interaction potential is expected to be much small, we could only keep the leading-order modification in the calculation. To be explicit, we neglect all the $\mathcal{O}(\delta^{2})$ terms, where the relative deviation $\delta$ is defined by $\delta\equiv(\widetilde{m}_{a}^{2}-{m}_{a}^{2})/{m}_{a}^{2}$. Given the quantum state of the boson cloud and the parameters of the gravitational atom, we can solve the frequency eigenvalue of the cloud and find out the rate of superradiant instability for the self-interacting boson.
@@
This condition ensures that the modification to the boson mass due to self-interaction is negligible compared to the particle’s original mass. Therefore, although the continued fraction equation generally provides a solution, the condition (30) must hold for the bound-state cloud to ensure the solution is physical. At the time the fast-spinning BH is born, for example when two primary BHs merge to a new BH or when a supernova explodes, the occupation number of bound state bosons is negligible. At the beginning of boson cloud growth, the small particle number keeps that the condition (30) is valid all through, while the self-interaction effect of ultralight boson is also very insignificant. When the bound state bosons increase by extracting the angular momentum of BH and eventually form a macroscopic cloud, the self-interaction term and the modification on the boson mass would become more and more important in the superradiance process. As we can find out, the relative deviation $\delta\propto N$.
@@ III The superradiance rate of the self-interacting boson cloud
In this section, we numerically compute the growth rate of self-interacting boson cloud. We utilize the methods introduced in the previous section to model the self-interaction effects of the boson cloud. In applying the continued fraction method, it is important to retain a sufficient number of modes to ensure sufficient numerical precision [^3] Previous studies [14, 70] have indicated that for a moderately small $\alpha$, the required mode number is on the order of a few hundred. In contrast, for very small $\alpha$, the required mode number increases substantially, exceeding one thousand. To assess the relative errors, we use the result corresponding to a mode number of approximately 5000 as the reference point.. For the black hole–boson cloud system parameters in our calculations, we choose a fixed BH mass, such that the gravitational fine-structure constant $\alpha$ is directly determined by the boson mass. We compute the superradiance rate for the self-interacting boson and compare it with that for the free boson.
@@ III.1 Different spins of BH
The spin of BH is the initial factor that trigger the superradiance process of the boson field. As the BH’s angular momentum is transferred to the bound state during the formation of the boson cloud, the BH spin gradually decreases until the superradiance condition is no longer satisfied. When a new BH forms through binary BH mergers or supernova explosions, it is generally expected to possess a high initial angular momentum. Additionally, the BH spin can increase via astrophysical processes, such as disk accretion. The dimensionless spin of astrophysical BHs can reach values exceeding 0.998 during their evolution [74, 75, 76, 77, 78, 79]. We calculate the growth rate of scalar boson—–defined as the imaginary part of the frequency eigenvalue—–for a range of BH dimensionless spins $\chi=0.9,0.95,0.99$. For comparison between the self-interacting and free bosons, we plot the free boson results using solid lines, while the growth rates accounting for boson self-interaction are shown with dashed lines, as illustrated in Fig. 1.
@@
In Fig. 1, we choose a moderate value of self-interaction strength. The energy scale of self-interaction $f_{a}$ is around $2.5\times 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$. We use the same $f_{a}$ in the calculation for self-interacting boson clouds. As for the gravitational atom, although the BH spins we consider range from 0.9 to 0.99, the spin difference of the BH from the initial stage when the boson cloud formation begins remains unchanged. This approach ensures that the boson cloud maintains an almost equal occupation number. We specify the benchmark parameters on the top of each figure.
From the results of Fig. 1, we find that for the self-interacting boson cloud the growth rates become smaller owing to the existence of self-interaction effect. The growth rates decrease by a half to one order of magnitude compared to the calculation results without self-interaction. In the free scalar field case, every solid line reaches a peak and then plunges eventually as the $\alpha$ increase. For the self-interacting boson, the results have similar behavior. Moreover, the peak positions for each BH spin remain approximately unchanged. That means in the case with self-interaction the superradiance condition (1) could be still applied.
@@ IV.1 GW strain
In this subsection, we focus on the strength of the GW signals from the pair annihilation process in the boson cloud. We are interested in the GW radiation process after the cloud saturation because the boson occupation number has reached the maximum and the GWs have also reached its peak emission power. The GW radiation process could last for a much longer timescale than the cloud formation. At that period, the cloud depletes gradually if only considering the annihilation to GWs and the occupation number of $\,\left|211\right\rangle$ state evolves as
@@
where $\tau_{\text{DC}}$ is the cloud depletion timescale through annihilation and is longer than $\tau_{\text{SR}}$ by a few of orders. Generally the altered maximum occupation number of bound state boson is smaller in the parameter space of interest when introducing the self-interaction, as can be seen from Sec. III.3. If considering a stronger coupling (smaller $f_{a}$), the maximum number of $N$ is suppressed more significant. During the depletion of cloud, the GW radiation power is proportional to $N^{2}(t)$, so the GW amplitude the detector receives on Earth, $h(t)$, is proportional to $N(t)$. On account of the scalar self-interaction, the maximum occupation number for the newly-formed boson cloud is $\widetilde{N}_{c}$. Thus, the peak GW amplitude is suppressed by the factor $\widetilde{N}_{c}/{N}_{c}$ (cf. Sec. III.3). For the boson cloud at a given age, the GW strain measured on Earth is also smaller than expected in the case of no self-interaction. In Sec. III, we has proposed that $\widetilde{N}_{c}$ is usually much smaller than ${N}_{c}$. In that respect, whether we consider the self-interaction effect could alter the GW amplitude significantly due to the change of cloud occupation number. Moreover, if the boson cloud has evolved with the age much longer than $\tau_{\text{DC}}$, the initial maximum occupation number $N(t=0)$ (which is equal to $\widetilde{N}_{c}$) becomes irrelevant and the GW amplitude $h$ roughly only depends on the annihilation rate to graviton $\Gamma_{a}$ and the time $t$. In this case, the concrete value of $f_{a}$ seems not so crucial. However, since the GW amplitude $h$ has become very small, this circumstance is out of interest.

### v2

@@ II.3 Self-interacting scalar bosons
Axions, or axion-like particles, are promising dark matter candidates. The searches for ultralight scalars like axions to date have not adequately accounted for self-interactions among particles in the superradiance process. Self-interactions, that could significantly alter superradiance dynamics, are generically present for many scalar models, including axion-like dark matter candidates and string axions. Besides the QCD axion arising as the result of the Peccei-Quinn symmetry broken at a high energy scale $f_{a}$, more general axion-like particles can be present in particular theory and are typically predicted to arise in string theory, in which the symmetry-breaking scale $f_{a}$ can fall in a wide mass range [9, 37]. From a particle physics perspective, the simplest interaction of a scalar field is the quartic self-interaction.
@@
Let us still begin with the scalar field Lagrangian (cf. Eq. (1)) on the Kerr metric. Self-interacting ultralight bosons generally arising through non-perturbative theory at some high energy scale give an effective potential. Any type of potential of the scalar field $\phi$ can be expanded to the lowest order with the mass term and the quartic self-interaction term. The strength of the mass term in the potential is proportional to the square of scalar mass, and the quartic self-interaction term proportional to the dimensionless self-interaction parameter $\lambda$. That is,
@@
$\lambda\approx 10^{-82}\left(\frac{m_{a}}{10^{-13}\penalty 10000\ \mathrm{eV}}\right)^{2}\left(\frac{10^{19}\penalty 10000\ \mathrm{GeV}}{f_{a}}\right)^{2}.$ (24)
@@
To the lowest order of $\phi$, heuristically Eq. (25) can be considered as the new Klein-Gordon equation on the Kerr metric and the deviation comes from the modification of the scalar mass. By considering the $\phi^{2}$ part as the background field, we can investigate this mass modification. The main contribution stems from the dominating energy level of bound state which grows at the maximum speed and occupies much more than others. The particle occupation number of the fastest growing level $N$ is proportional to the square of its field amplitude ($N\propto|\phi|^{2}$). We separate this part individually and define $\Pi$ to quantify the contribution of bound state wave function except $N$, i.e. $N\Pi={\phi^{2}}/\left.{f_{a}^{2}}\right$.. The quantity $\Pi$ is only determined by the distribution of the wave function and the scale of $f_{a}$. We are going to evaluate this modification in a more realistic situation for the superradiant growth. The realistic approach is based on the fact that for the gravitational atom’s parameter space of interest the growth timescale of the scalar cloud is much longer than the oscillation timescale of the particles, ${\omega}^{-1}$ ($\approx m_{a}^{-1}$). Also, the cloud growth timescale is much slower than the timescale of the particles’ orbital motion in the vicinity of the BH. For these reasons, we can perform the time- and spatial averaging when investigating $\Pi$ within one period and calculating the scalar mass deviation. Using similar notation as in Eq. (10) and Eq. (11), we can then calculate $\Pi$ with
@@
where $\Gamma_{\rm SR}$ is the growth rate of bosons, $t$ is the time and the bars on the symbols denote the average over spatial variables. In the non-relativistic limit, we already know the analytical expressions of radial and angular functions. The radial dependence of function, $R$, becomes radial hydrogen wave function, while the angular dependence of function, $S$, becomes the spherical harmonics. Based on the more thorough calculation, although the analytical results provide a good approximation, we carry out the estimation within the full Kerr background. We can investigate the mean values of the functions around the cloud region. Note that the radial and angular functions both obey the normalization condition. As for the occupation number at a specific time, we can estimate $N$ through the spin difference of the BH due to the conservation of angular momentum. When only considering the fastest growing level of bound state, the scalar cloud is dominated by the particles of single state and the extracted angular momentum from Kerr BH is stored in the same state. The occupation number of the fastest growing state is
@@
where $\varDelta\chi\lesssim\mathcal{O}(0.1)$. Owing to the hierarchy of growth rates for different superradiant states, the newly-formed BH is always dominated by one bound state of scalar boson, and higher states would take much longer time to grow to the saturation. Thus, we consider almost all bound-state bosons occupy the same energy level here.
@@ II.4 The spectra of the gravitational atom
We can apply the modification on the recurrence relation Eq. (17) and obtain the new continued fraction equation for self-interacting bound state boson. While the deviation of boson mass induced by the self-interaction potential is expected to be much small, we could only keep the leading-order modification in the calculation. To be explicit, we neglect all the $\mathcal{O}(\delta^{2})$ terms, where the relative deviation $\delta$ is defined by $\delta\equiv(\widetilde{m}_{a}^{2}-{m}_{a}^{2})/{m}_{a}^{2}$. Given the quantum state of the boson cloud and the parameters of the gravitational atom, we can solve the frequency eigenvalue of the cloud and find out the rate of superradiant instability for the self-interacting boson. In general, modifications to the energy spectra of a gravitational atom can be classified into two distinct types. First, changes of the imaginary part of the complex eigenfrequency directly affect the rate of superradiant instability in the self-interacting boson cloud, governing the exponential growth of the bound-state occupation number. As we will show later, the self-interaction could potentially suppress the growth rate of boson cloud; and this effect will be discussed in detail in Sec. III. Second, the modification to the real part of the eigenfrequency induces energy-level shift of the boson cloud, which will be specifically analyzed in Sec. IV.
@@
This condition ensures that the modification to the boson mass due to self-interaction is negligible compared to the particle’s original mass. Therefore, although the continued fraction equation generally provides a solution, the condition (30) must hold for the bound-state cloud to ensure the solution is physical. At the time the fast-spinning BH is born, for example when two primary BHs merge to a new BH or when a supernova explodes, the occupation number of bound state bosons is negligible. At the beginning of boson cloud growth, the small particle number keeps that the condition (30) is valid all through, while the self-interaction effect of ultralight boson is also very insignificant. When the bound-state bosons are produced by extracting the angular momentum of BH and eventually form a macroscopic cloud, the self-interaction term and the modification on the boson mass would become more and more important in the superradiance process. As we can find out, the relative deviation $\delta\propto N$.
It is worth to have more discussion on the conditions required to ensure the validity of the numerical methods introduced in this section. Firstly, we adopt the continued fraction method to solve both real and imaginary components of the boson eigenfrequency. The continued fraction equation is derived based on the Klein-Gordon equation on Kerr metric, so it is applicable for different value of $\alpha$. Both the superradiant growth rate and the energy level of the boson cloud can be calculated numerically for the general cases in which $\alpha$ might be relatively large. Secondly, the treatment of the self-interaction term as the modification to the boson mass is valid only when the occupation number satisfies $\frac{1}{3!}N\Pi\ll 1$. As the occupation number grows exponentially during the boson cloud formation, perhaps the condition may no longer satisfy once the occupation number is sufficiently large. However, the self-interaction effects generally bring a new upper limit that restrict the growth of occupation number. When the occupation number approaches to the new bound, the growth rate approaches to zero. With more detailed calculations, we find that for the parameter space of interest, the boson cloud growth always ceases much earlier before the condition (30) breaks down. That indicates it is safe to treat the self-interaction as the perturbative modification on boson mass when considering the newly-formed self-interacting boson cloud. Moreover, as for the drastic bosenova effect [37, 72, 73], current studies [72, 73] had performed the numerical simulation for the axion cloud growth and the results had indicated during the cloud growth, the strong self-interactions would lead to the nonperturbative collapse of the cloud eventually. However, the upper bound of cloud occupation number also potentially prevents the bosenova collapse, cause the bosenova occurs only when the cloud occupation number exceeds the critical amount. If the new upper limit of occupation number is small than the critical value needed to trigger bosenova, there is low probability for the bosenova to occur during the boson cloud formation. Nonetheless, the numerical approaches introduced in this section assume the whole boson cloud is dominated by the fastest growing state, and neglect the contribution from other higher states. Ref. [51] had proposed the moderate and large self-interaction can transfer the bosons to other higher states and several bound states are excited in the boson cloud. This work shows that when considering the energy exchange and level mixing between different bound states of the gravitational atom, the quartic self-interaction will cause the boson cloud to saturate at a new multi-state equilibrium, while the original fastest-growing state will have a lower occupation number. Although the growth timescales of higher superradiant states are much longer than the fastest one, the self-interaction can excite the transition between these states. After other higher states are populated, they bring extra effects on the growth of the dominant state. Yet these effects are far beyond the scope of our work. In the following calculations, we omit the contribution from the multiple states except the self-interaction of the fastest growing state itself.
@@ III The superradiance rate of the self-interacting boson cloud
In this section, we numerically compute the growth rate of self-interacting boson cloud. We utilize the methods introduced in the previous section to model the self-interaction effects of the boson cloud. In applying the continued fraction method, it is important to retain a sufficient number of modes to ensure sufficient numerical precision [^3] Previous studies [15, 75] have indicated that for a moderately small $\alpha$, the required mode number is on the order of a few hundred. In contrast, for very small $\alpha$, the required mode number increases substantially, exceeding one thousand. . For the BH–boson cloud system parameters in our calculations, we choose a fixed BH mass, such that the gravitational fine-structure constant $\alpha$ is directly determined by the boson mass. We compute the superradiance rate for the self-interacting boson and compare it with that for the free boson. Due to the significant disparity in growth rates across different superradiant states, the newly-formed BH is always dominated by single bound state of scalar boson. Higher-order states exhibit substantially longer saturation timescales. Thus, we model the system with the approximation that virtually all bound-state bosons populate the single energy level which has the fastest growth rate, and neglect potential level mixing [51]. For the parameter space of interest, we consider only $\left|211\right\rangle$ state in this section.
@@ III.1 Different spins of BH
The spin of BH is the initial factor that triggers the superradiance process of the boson field. As the BH’s angular momentum is transferred to the bound state during the formation of the boson cloud, the BH spin gradually decreases until the superradiance condition is no longer satisfied. When a new BH forms through binary BH mergers or supernova explosions, it is generally expected to possess a high initial angular momentum. Additionally, the BH spin can increase via astrophysical processes, such as disk accretion. The dimensionless spin of astrophysical BHs can reach values exceeding 0.998 during their evolution [79, 80, 81, 82, 83, 84]. We calculate the growth rate of scalar boson—–defined as the imaginary part of the frequency eigenvalue—–for a range of BH dimensionless spins $\chi=0.9,0.95,0.99$. For the comparison between the self-interacting bosons and the free bosons, we plot the free boson results using solid lines, while the growth rates accounting for boson self-interaction are shown with dashed lines, as illustrated in Fig. 1.
@@
In Fig. 1, we choose a moderate value of self-interaction strength. The energy scale of self-interaction $f_{a}$ is around $2.5\times 10^{18}\penalty 10000\ \mathrm{GeV}$. We use the same $f_{a}$ in the calculation for self-interacting boson clouds. As for the gravitational atom, although the BH spins we consider range from 0.9 to 0.99, the spin difference of the BH from the initial stage when the boson cloud formation begins remains unchanged. This approach ensures that the boson cloud maintains an almost equal occupation number. We specify the benchmark parameters on the top of each figure.
From the results of Fig. 1, we find that for the self-interacting boson cloud the growth rates become smaller owing to the existence of self-interaction effect. The growth rates decrease by a half to one order of magnitude compared to the calculation results without self-interaction. In the free scalar field case, every solid line reaches a peak and then drops eventually as the $\alpha$ increases. For the self-interacting boson, the results have similar behavior. We can see that around these highest points, with the increase of $\alpha$, the growth rates decline and approach to zero. These critical values of $\alpha$ correspond to the superradiance condition for different BH spin. The allowed range of $\alpha$ for the superradiant growth of boson cloud is dependent on the value of $\chi$. For the larger BH spin, the boson cloud can grow via the superradiance process for the larger $\alpha$. These results can be understood from the superradiance condition (4). Moreover, when considering the self-interaction effects the upper bounds of the $\alpha$ ranges for each BH spin remain approximately unchanged. That means in the case with self-interaction the superradiance condition would be still applicable.
@@ III.3 Different occupation numbers of boson cloud
Before moving to the next section, we calculate the relative errors when solving for the superradiance rates of the self-interacting boson clouds. The relative error partly comes from the truncation we choose when solving the transcendental equation (20). We analyze the number of the terms in the continued fraction required for the convergence of solution. We define the relative error criterion to determine the convergence of eigenfrequency as
Figure 4: The imaginary part of eigenfrequency obtained from the numerical method, for different numbers of terms kept in continued fraction. For the smaller $\alpha$, the solution converges more slowly and more terms in the continued fraction are required. The results are for the benchmark parameters, but similar conclusion holds for a wide range of parameter space.
$\epsilon_{n}\equiv\left|\frac{\omega^{(n)}-\omega^{(n-50)}}{\omega^{(n)}}\right|<10^{-4},$ (32)
where $\omega^{(k)}$ denotes the imaginary part of frequency eigenvalue computed with $k$ fraction terms. We show the result of imaginary eigenfrequency with different mode number in Fig. 4. It can be seen from Fig. 4 that it takes more terms to be kept to ensure the convergence of the solution for a smaller value of $\alpha$. In the calculations of this section, we choose the term number $\sim 1000$ to ensure the convergence while optimizing computational efficiency.
@@ IV.1 GW strain
In this subsection, we focus on the strength of the GW signals from the pair annihilation process in the boson cloud. We are interested in the GW radiation process after the cloud saturation because the boson occupation number has reached the maximum and the GWs have also reached its peak emission power. The GW radiation process could last for a much longer timescale than the cloud formation. During that period, the cloud depletes gradually if only considering the annihilation of bosons to GWs and the occupation number of $\left|211\right\rangle$ state evolves as
@@
where $\tau_{\text{DC}}=(\Gamma_{a}N(t=0))^{-1}$ is the cloud depletion timescale through annihilation and is longer than $\tau_{\text{SR}}$ by a few of orders. $\Gamma_{a}$ is the annihilation rate of bosons, and $t=0$ corresponds to the onset of boson cloud saturation. Generally, the altered maximum occupation number of bound state boson is smaller than ${N}_{c}$ in the parameter space of interest when introducing the self-interaction, as shown in Sec. III.3. If considering a stronger coupling (smaller $f_{a}$), the maximum number of $N$ is suppressed more significantly. In the depletion process of cloud, the GW radiation power is proportional to $N^{2}(t)$, so the GW amplitude received by the detector $h(t)$, is proportional to $N(t)$. On account of the scalar self-interaction, the maximum occupation number for the newly-formed boson cloud is $\widetilde{N}_{c}$. Thus, the peak GW amplitude is suppressed by the factor $\widetilde{N}_{c}/{N}_{c}$ (cf. Sec. III.3). In Sec. III, we has proposed that $\widetilde{N}_{c}$ is usually much smaller than ${N}_{c}$. In that respect, whether we consider the self-interaction effect could alter the peak GW amplitude significantly due to the change of cloud occupation number.
Figure 5: The suppression factor of peak GW amplitude for the continuous GWs emitted by the self-interacting boson cloud.
In Fig. 5, we demonstrate the GW amplitude suppression factor $\widetilde{N}_{c}/{N}_{c}$ as the function of the self-interaction scale $f_{a}$. It can be seen from the plot that, for the smaller self-interaction scale, the maximum occupation number of boson cloud would decrease significantly, leading to a lower peak GW strain $h$. Moreover, because of the GW radiation, the boson cloud will progressively lose its energy after its saturation and its occupation number will decrease with time [^4] This is always correct no matter whether the self-interaction is included. The difference when introducing the self-interaction effect is the alteration of occupation number decreasing rate.. In contrast to the newly-formed boson cloud, for the boson cloud at a given age, the emitted GW strain will also become smaller with time. It can be seen from Eq. (33) that, if the boson cloud has evolved with the age much longer than $\tau_{\text{DC}}$, its occupation number could be estimated as $N(t)\sim{\tau_{\text{DC}}N(t=0)}/{t}=(\Gamma_{a}t)^{-1}$. Therefore, when the age of boson cloud is much longer than $\tau_{\text{DC}}$, the initial maximum occupation number $N(t=0)$ (which is equal to $\widetilde{N}_{c}$) becomes irrelevant and the GW amplitude $h$ roughly only depends on the annihilation rate to graviton $\Gamma_{a}$ and the time $t$. In this case, the concrete value of $f_{a}$ seems not so crucial. However, since the GW amplitude $h$ has become very small, this circumstance is out of interest.